OpenStudy (anonymous):

sqrt 48x^3 y^8 z^3 <--- must be in simplest radical form , show work please

OpenStudy (anonymous):

4y^4 x z sqrtxz3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[\sqrt{48 x^2 y^8 z^3}\] \[\sqrt{48\text{ }}\sqrt{x^2}\sqrt{y^8 }\sqrt{z^3\text{ }}\] \[4\sqrt{3\text{ }}x y^4 \sqrt{z^3\text{ }}\]

OpenStudy (anonymous):

thank you